OpenStudy (manny_9):

If y varies directly as x, and y = 16 as x = 4, find y for the x-value of 2. 5 6 7 8

Directrix (directrix):

y=k*x 16 = kx k = 4 y=4x y = 4 *2 y = 8 Check that.

Directrix (directrix):

@manny_9

OpenStudy (manny_9):

i got the same answer. thanks

Directrix (directrix):

You're welcome.